# Preserving Privacy Without Compromising Accuracy: Machine Unlearning for Handwritten Text Recognition

### v1

@@ Abstract
Handwritten Text Recognition (HTR) is essential for document analysis and digitization. However, handwritten data often contains user-identifiable information, such as unique handwriting styles and personal lexicon choices, which can compromise privacy and erode trust in AI services. Legislation like the “right to be forgotten” underscores the necessity for methods that can expunge sensitive information from trained models. Machine unlearning addresses this by selectively removing specific data from models without necessitating complete retraining. Yet, it frequently encounters a privacy-accuracy tradeoff, where safeguarding privacy leads to diminished model performance. In this paper, we introduce a novel two-stage unlearning strategy for a multi-head transformer-based HTR model, integrating pruning and random labeling. Our proposed method utilizes a writer classification head both as an indicator and a trigger for unlearning, while maintaining the efficacy of the recognition head. To our knowledge, this represents the first comprehensive exploration of machine unlearning within HTR tasks. We further employ Membership Inference Attacks (MIA) to evaluate the effectiveness of unlearning user-identifiable information. Extensive experiments demonstrate that our approach effectively preserves privacy while maintaining model accuracy, paving the way for new research directions in the document analysis community. Our code will be publicly available upon acceptance.
@@ 1 Introduction
In this paper, we propose a novel approach to enable efficient machine unlearning in HTR systems without compromising accuracy. We introduce an encoder-only transformer-based model as a baseline for HTR tasks, enhanced with a handwriting style classification head. This two-task model not only serves as an indicator of how much style information (the user-identifiable component) is memorized during training but also provides a lever to do machine unlearning. By applying neural pruning on the properly trained model and random label assignment to the style classification head, we effectively unlearn the user-identifiable information from the model. We further utilize membership inference attacks (MIA) [19] to evaluate whether user-identifiable information is eliminated as users’ request, demonstrating the effectiveness of our method.
@@
[RQ3] Can employing random labeling in the writer classification head effectively eliminate user-identifiable information on request, without harming the recognition head’s performance?
@@ 3.2 Solution Formulation
Stage II: Random Labeling. The forget set $D_{\text{forget}}$ is further unlearned by introducing data with random writer IDs.
@@ 3.4 Multi-head Baseline Model
In the single-head framework, machine unlearning techniques such as random labeling can only be applied through the recognition head, which inevitably leads to a decline in recognition performance on the forget set. To overcome this limitation, we enhance the single-head architecture by incorporating a special $[CLS]$ token alongside the handwritten text image $X$ as input and introducing a writer classification head $H_{w}$ in addition to the recognition head $H_{r}$. This modification is formulated as follows: $F_{c}^{{}^{\prime}}=M_{cnn}([CLS],X)$, $F_{t}^{{}^{\prime}}=M_{tran}(F_{c}^{{}^{\prime}})$, with the recognition head predicting the text $Y=H_{r}(F_{t}^{{}^{\prime}})$ and the writer classification head producing the writer-specific identity $Id=H_{w}(F_{t}^{{}^{\prime}})$. By capturing user-specific attributes, such as handwriting style and inherent lexicon preferences, this design enables the model to associate such features with a unique user identity. Consequently, the writer classification head serves as an indicator of the extent to which the model retains user-specific information.
@@ 3.5 Unlearning Stage I: Neural Pruning
3.6 Unlearning Stage II: Random Labeling
The knowledge associated with both the forget and retain sets is deeply coupled across all neurons in the model $M$. Consequently, the first stage of neural pruning, which removes only the dominant neurons sensitive to the forget set, serves merely as an initial step. At this point, the pruned model $M^{*}$ still retains some user-identifiable information that needs to be forgotten, and there is a certain level of information decline affecting the retain set. Consequently, targeted unlearning techniques are required to enhance the retain set while ensuring the complete removal of user-identifiable information from the forget set.
We implement a random labeling strategy that only applies random labels to the writer classification head, leaving the recognition head unchanged. The writer ID corresponds to all handwriting samples produced by a single writer. Hence, writer classification can be seen as identifying user-specific information, represented as handwriting style for the visual modality and inherent lexicon usage for the language modality.
To achieve this, we randomly reassign labels to the samples in $D_{\text{forget}}$, ensuring that the reassigned labels are different from the original ones, thereby creating $D_{\text{forget}}^{\prime}$. Next, we merge $D_{\text{forget}}^{\prime}$ with $D_{\text{retain}}$ to obtain the new training dataset $D_{\text{train}}^{\prime}$ and use it to update $M^{*}$, ultimately yielding the unlearned model $M^{\prime}$.
@@ 4.2 Dataset and Metrics
Here, $S$, $I$ and $D$ represent the number of substitutions, insertions, and deletions, respectively, required to transform one string into the other, either at the character or word level. $N$ denotes the total number of characters in the ground truth for CER and the total number of words in the ground truth for WER. A lower CER or WER indicates better HTR performance with fewer recognition errors.
@@ 4.5.7 Neural Pruning Analysis
After completing all pruning experiments across each module, we can address research question [RQ2] by concluding that neural pruning can partially remove user-identifiable information. However, due to the deep coupling of knowledge within the model for both the forget set and the retain set, encompassing both user-identifiable and recognition-related information, pruning alone is insufficient to fully eliminate user-identifiable information while retaining the model’s useful knowledge. Consequently, an additional step is necessary to perform direct unlearning.
@@
4.6 Random Labeling Experiments
From the analysis in the previous section, we select our pruned model $M^{*}$ with a pruning rate of 40% for the embedding module, and 20%, 40%, and 20% for the shallow, middle, and deep blocks of the self-attention modules, respectively. Additionally, a pruning rate of 20% is applied to all feed-forward modules. Both the baseline model $M$ and the pruned model $M^{*}$ are then fine-tuned on the updated training set $D_{\text{train}}^{\prime}$, which comprises the randomly labeled forget set $D_{\text{forget}}^{\prime}$, where the user IDs in the forget set are replaced with random user IDs, excluding the real ones, and the retain set $D_{\text{retain}}$. The results are shown for different iterations as detailed in Tab. 10.
@@ 4.6 Random Labeling Experiments
4.6.1 Membership Inference Evaluation
As previously discussed, the writer classification head serves as an indicator. However, achieving 0% accuracy does not necessarily confirm non-membership for the forget set. To address this, we conduct a membership inference evaluation, as shown in Tab. 11. For both the baseline and pruned models, with and without random labeling, members (retain set) are classified as “seen” with over 75% accuracy, while non-members (test set) are classified as “unseen” with an accuracy range of 51% to 59%. These outcomes align with expectations: members have a high classification probability as they were seen during training, while non-members approach the random guessing baseline (50%), reflecting the model’s lack of prior knowledge about them.
Thus, we can now address research question [RQ3] by concluding that using random labeling in the writer classification head effectively removes user-identifiable information without negatively impacting recognition performance.
Table 11: Membership inference evaluation with random labeling applied to both the baseline and pruned models.

### v2

@@ Abstract
Handwritten Text Recognition (HTR) is crucial for document digitization, but handwritten data can contain user-identifiable features, like unique writing styles, posing privacy risks. Regulations such as the “right to be forgotten” require models to remove these sensitive traces without full retraining. We introduce a practical encoder-only transformer baseline as a robust reference for future HTR research. Building on this, we propose a two-stage unlearning framework for multihead transformer HTR models. Our method combines neural pruning with machine unlearning applied to a writer classification head, ensuring sensitive information is removed while preserving the recognition head. We also present Writer-ID Confusion (WIC), a method that forces the forget set to follow a uniform distribution over writer identities, unlearning user-specific cues while maintaining text recognition performance. We compare WIC to Random Labeling, Fisher Forgetting, Amnesiac Unlearning, and DELETE within our prune-unlearn pipeline and consistently achieve better privacy and accuracy trade-offs. This is the first systematic study of machine unlearning for HTR. Using metrics such as Accuracy, Character Error Rate (CER), Word Error Rate (WER), and Membership Inference Attacks (MIA) on the IAM and CVL datasets, we demonstrate that our method achieves state-of-the-art or superior performance for effective unlearning. These experiments show that our approach effectively safeguards privacy without compromising accuracy, opening new directions for document analysis research. Our code is publicly available at https://github.com/leitro/WIC-WriterIDConfusion-MachineUnlearning.Preprint version of the accepted manuscript in Pattern Recognition, available at DOI: https://doi.org/10.1016/j.patcog.2025.112411
@@ 1 Introduction
In this paper, we propose a novel approach to enable efficient machine unlearning in HTR systems without compromising accuracy. We introduce an encoder-only transformer-based model as a baseline for HTR tasks, enhanced with a handwriting style classification head. This two-task model not only serves as an indicator of how much style information (the user-identifiable component) is memorized during training but also provides a lever to do machine unlearning. By applying neural pruning on the properly trained model and subsequent unlearning techniques targeting the style classification head, we effectively remove user-identifiable information from the model. Additionally, we propose a Writer-ID Confusion (WIC) method that promotes uncertainty by enforcing a uniform distribution over writer identity predictions for the forget set, while preserving cross-entropy loss for the retain set. We further utilize membership inference attacks (MIA) [19] to evaluate whether user-identifiable information is eliminated as users request, demonstrating the effectiveness of our method.
@@
[RQ3] Can employing machine unlearning in the writer classification head effectively eliminate user-identifiable information on request, without harming the recognition head’s performance?
@@
    We propose a novel Writer-ID Confusion (WIC) method that facilitates unlearning of user-identifiable information by enforcing a uniform distribution over writer ID predictions for the forget set, while retaining standard cross-entropy training on the retain set. This introduces uncertainty in identity-specific predictions, aiding effective machine unlearning.
  • 5.
@@ 3.2 Solution Formulation
Stage II: Machine Unlearning. The forget set $D_{\text{forget}}$ is further processed using machine unlearning techniques, including our proposed WIC method.
@@ 3.4 Multi-head Baseline Model
In the single-head framework, machine unlearning techniques such as random labeling can only be applied through the recognition head, which inevitably leads to a decline in recognition performance on the forget set. To overcome this limitation, we enhance the single-head architecture by incorporating a special $[CLS]$ token alongside the handwritten text image $X$ as input and introducing a writer classification head $H_{w}$ in addition to the recognition head $H_{r}$. This modification is formulated as follows: $F_{c}^{{}^{\prime}}=M_{cnn}([CLS],X)$, $F_{t}^{{}^{\prime}}=M_{tran}(F_{c}^{{}^{\prime}})$, with the recognition head predicting the text $Y_{ocr}=H_{r}(F_{t}^{{}^{\prime}})$ and the writer classification head producing the writer-specific identity $Y_{wid}=H_{w}(F_{t}^{{}^{\prime}})$. By capturing user-specific attributes, such as handwriting style and inherent lexicon preferences, this design enables the model to associate such features with a unique user identity. Consequently, the writer classification head serves as an indicator of the extent to which the model retains user-specific information.
@@ 3.5 Unlearning Stage I: Neural Pruning
3.6 Unlearning Stage II: Post‑hoc Machine Unlearning
After neural pruning in Stage I, we obtain a sparsified model $M^{*}$ that has removed the most writer-sensitive neurons. In Stage II, we further eliminate residual writer-identity information while preserving the accuracy of character-level recognition. To achieve this, we investigate four state-of-the-art unlearning methods including Random Labeling, Fisher Forgetting, Amnesiac Unlearning, and DELETE, and propose a novel Writer-ID Confusion (WIC) loss. All methods operate on mini-batches $\mathcal{B}=\mathcal{B}_{\mathrm{retain}}\cup\mathcal{B}_{\mathrm{forget}}$, with retain and forget samples defined in Sec. 3.1. We denote by $\hat{\mathbf{y}}_{\mathrm{wid}}$ the logits from the writer classification head $H_{w}$, and by $\mathcal{L}_{\text{ocr}}$ the sequence-recognition loss from the baseline HTR model.
3.6.1 Random Labeling
We selectively corrupt the writer labels only for the forget set, transforming $D_{\text{forget}}$ into $D_{\text{forget}}^{\prime}$ through a one-to-one random permutation $perm:W_{\text{forget}}\rightarrow W\setminus W_{\text{forget}}$. The resulting training objective becomes:
$\mathcal{L}_{\text{Rand}}=\text{CE}\bigl(\hat{\mathbf{y}}_{\text{wid}},\mathbf{y}_{\text{perm}}\bigr)+\beta\mathcal{L}_{\text{ocr}},$ (2)
where $\mathbf{y}_{\text{perm}}$ coincides with the original labels for the retain set and the randomly permuted labels for the forget set. This approach deliberately forces misclassification of forgotten writers while minimally impacting recognition performance.
3.6.2 Fisher Forgetting
Fisher Forgetting [32] is a method that selectively removes information about specific data points from a trained neural network. It applies a single Newton step using the Fisher Information Matrix (FIM), along with geometry-matched Gaussian noise, allowing efficient unlearning without retraining. The original formulation is as follows:
$\theta_{\text{scrub}}=\theta-(F+\mu I)^{-1}g_{f}+\mathcal{N}\left(0,\tau^{2}(F+\mu I)^{-1}\right)$ (3)
In this equation, $\theta$ represents the original model parameters, $F$ is the Fisher Information Matrix estimated on the forget set, $\mu$ is a small positive damping term, $g_{f}$ is the gradient of the loss with respect to the forget set, and $\mathcal{N}(\cdot)$ denotes Gaussian noise with covariance scaled by the noise parameter $\tau^{2}$.
However, this formulation is tailored for image classification tasks such as MNIST and CIFAR10. For our sequence recognition task, we adapt the equation to a simpler form that operates on the diagonal of the Fisher Information Matrix:
$\theta_{\text{scrub}}=\theta-\alpha\frac{g_{f}}{F+\mu}$ (4)
Here, $F$ denotes the diagonal Fisher Information Matrix computed on the forget set, and $\alpha$ is the update step size (set to 0.01).
3.6.3 Amnesiac Unlearning
Amnesiac Unlearning [33] aims to reverse the training dynamics for forgotten data by iteratively applying scaled negative gradient steps:
$\theta^{(t+1)}=\theta^{(t)}+\alpha\,\nabla_{\theta}\,\mathcal{L}\,\bigl(x_{f}^{t},y_{f}^{t}\bigr),$ (5)
where $(x_{f}^{t},y_{f}^{t})$ is the $t$-th mini-batch drawn from $D_{\text{forget}}$, $\nabla_{\theta}$ is the gradient operator with respect to the model parameters $\theta$, and $\alpha$ is a small scaling factor (approximately $10^{-6}$). By reversing parameter updates, the method systematically erases previously learned information related to forgotten data.
3.6.4 DELETE Distillation
DELETE [34] removes sensitive classes by pairing a frozen teacher $M^{T}$ with a trainable student $M^{S}$. For every mini‑batch, a decoupled distillation mask $\mathbf{m}\!\in\!\mathbb{R}^{N\times K}$ sets the logits of the classes that must be forgotten to -$\infty$, leaving only the logits of the retain classes unchanged. Applying softmax to the masked teacher outputs yields the teacher distribution $\mathbf{p}_{r}^{T}$ that is defined solely over the retain sub‑space, ensuring no gradient ever flows through the forgotten identities.
Training minimises a single, masked knowledge‑distillation objective
$\mathcal{L}_{\text{DELETE}}=\lambda_{\mathrm{KD}}\;\mathrm{KL}\!\bigl(\mathbf{p}_{r}^{T}\,\|\,\mathbf{p}_{r}^{S}\bigr),$ (6)
where $\mathbf{p}_{r}^{S}$ is the student softmax restricted to the same retain classes and $\lambda_{\mathrm{KD}}$ is a balancing coefficient (set to $1$ in our experiments). Because the teacher is frozen and the forget‑class logits are explicitly nulled, the student converges to a state in which it no longer encodes information about the forgotten identities while preserving performance on the retain set.
3.6.5 Fine‑Tuning
Fine-tuning (FT) is a widely used yet surprisingly strong baseline for machine unlearning. Upon the deletion request, all samples belonging to the forget classes are removed and the remaining data are used to continue training for a few additional epochs with a small learning rate. Because the procedure only sees retain data, the classifier naturally drifts away from the forgotten classes while preserving its useful knowledge.
Formally, given a mini‑batch of retain samples with one‑hot labels $\mathbf{y}_{r}$ and student logits $\mathbf{z}_{r}$ (restricted to the retain classes), FT minimises the standard cross‑entropy
$\mathcal{L}_{\text{FT}}\;=\;\mathrm{CE}\,\bigl(\hat{\mathbf{y}}_{r},\,\mathbf{y}_{r}\bigr)$ (7)
In our experiments, due to a budgeted number of unlearning iterations, and because the SOTA methods alone do not perform well, we first apply the SOTA methods and then use a fine-tuning strategy to achieve optimal performance.
3.6.6 Our Proposed Writer-ID Confusion (WIC)
To specifically target the removal of sensitive writer identity information without compromising the model’s handwritten text recognition performance, we propose the Writer-ID Confusion (WIC) method as shown in Fig. 3. Unlike traditional random labeling approaches, which explicitly assign incorrect writer labels, our method introduces uncertainty by enforcing a uniform distribution over writer identity predictions for samples belonging to the forget set.
Figure 3: Pipeline of our proposed Writer-ID Confusion (WIC) method, which enforces a uniform writer ID distribution on the forget set while applying standard cross-entropy loss on the retain set.
Formally, the WIC loss is defined as follows:
$\mathcal{L}_{\text{WIC}}=\frac{1}{|\mathcal{B}_{\text{retain}}|}\sum_{i\in\mathcal{B}_{\text{retain}}}\text{CE}(\hat{\mathbf{y}}_{\text{wid}}^{i},y_{\text{wid}}^{i})+\lambda\frac{1}{|\mathcal{B}_{\text{forget}}|}\sum_{j\in\mathcal{B}_{\text{forget}}}\text{KL}\left(\text{softmax}(\hat{\mathbf{y}}_{\text{wid}}^{j})\parallel\mathbf{u}\right)$ (8)
Here, $\hat{\mathbf{y}}_{\text{wid}}$ represents the output logits from the writer classification head $H_{w}$, and $y$ denotes the true writer labels. The set $\mathcal{B}_{\text{retain}}$ consists of indices for retain set samples, while $\mathcal{B}_{\text{forget}}$ corresponds to forget set samples within a given mini-batch. The term CE represents the standard cross-entropy loss applied exclusively to the retain set to preserve the correct identification of non-sensitive writers. The KL divergence term forces the model’s writer-ID predictions for forget set samples to approximate a uniform distribution, denoted by $\mathbf{u}$, effectively removing the ability to identify these writers.
The total training objective is thus formulated as:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{WIC}}+\beta\mathcal{L}_{\text{rec}}$ (9)
where $\mathcal{L}\_{\text{rec}}$ is the sequence-level recognition loss, ensuring the model maintains high accuracy for handwritten text transcription. The hyperparameter $\lambda$ balances the degree of writer-ID forgetting, while setting $\beta=1$ maintains the standard recognition performance.
@@ 4.2 Dataset and Metrics
Here, $S$, II and $D$ represent the number of substitutions, insertions, and deletions, respectively, required to transform one string into the other, either at the character or word level. $N$ denotes the total number of characters in the ground truth for CER and the total number of words in the ground truth for WER. A lower CER or WER indicates better HTR performance with fewer recognition errors.
@@ 4.5.7 Neural Pruning Analysis
The Neural Pruning stage is illustrated in Fig. 5. After completing all pruning experiments across each module, we can address research question [RQ2] by concluding that neural pruning can partially remove user-identifiable information. However, due to the deep coupling of knowledge within the model for both the forget set and the retain set, encompassing both user-identifiable and recognition-related information, pruning alone is insufficient to fully eliminate user-identifiable information while retaining the model’s useful knowledge. Consequently, an additional step is necessary to perform direct unlearning.
Figure 5: Stage I: Prune 40% of the embedding module (yellow); 20%, 40%, and 20% of the shallow, middle, and deep self-attention blocks (blue, red, gray), respectively; and 20% of all feed-forward modules (green, purple, cyan).
While our empirical determination of pruning rates and configurations aimed to systematically uncover patterns related to layer sensitivity and user-identifiable information, it was not intended to suggest repetitive manual tuning in future scenarios. Instead, our comprehensive analysis serves as a foundation for developing more principled, automated pruning strategies. The documented patterns from our study can guide future research, significantly streamlining the unlearning process in new contexts. Additionally, we recognize that neural pruning alone does not fully separate information between the forget and retain sets, as indicated by the concurrent increase in recognition errors. This deep coupling highlights the need for integrating neural pruning with subsequent stages, such as random labeling, to effectively remove user-specific information while preserving recognition accuracy, thereby underscoring their complementary roles in achieving desired privacy outcomes.
@@
4.6 Final Results using the prune-unlearn pipeline with sota MU methods
In this section, we evaluate the effectiveness of our prune-unlearn pipeline when integrated with several state-of-the-art machine unlearning techniques: Random Labeling, Fisher Forgetting, Amnesiac Unlearning, DELETE, and our proposed WIC method. We compare these approaches based on their ability to remove user-identifiable information while maintaining handwriting text recognition performance. To assess the level of unlearning, we perform membership inference attacks (MIA) on all methods. The goal is for the forget set to behave similarly to the test set, indicating non-membership, while the retain set continues to exhibit characteristics of training members.
For Random Labeling, we conduct more detailed experiments across different unlearning iterations. To ensure a fair and consistent comparison across all methods, we fix the total number of training iterations to 10,000 and use a batch size of 64 throughout.
4.6.1 Random Labeling Experiments
Both the baseline model $M$ and the pruned model $M^{*}$ are then fine-tuned on the updated training set $D_{\text{train}}^{\prime}$, which comprises the randomly labeled forget set $D_{\text{forget}}^{\prime}$, where the user IDs in the forget set are replaced with random user IDs, excluding the real ones, and the retain set $D_{\text{retain}}$. The results are shown for different iterations as detailed in Tab. 10.
@@ 4.6.1 Random Labeling Experiments
Membership Inference Evaluation
As previously discussed, the writer classification head serves as an indicator. However, achieving 0% accuracy does not necessarily confirm non-membership for the forget set. To address this, we conduct a membership inference evaluation, as shown in Tab. 11. The forget set is initially part of the member category in the baseline recognition model, since it is included in the training data. After applying unlearning, our goal is for the forget set to exhibit behavior similar to non-members, effectively transitioning to an unseen status in the unlearned model. For both the baseline and pruned models, with and without random labeling, members (retain set) are classified as “seen” with over 75% accuracy, while non-members (test set) are classified as “unseen” with an accuracy range of 51% to 59%. These outcomes align with expectations: members have a high classification probability as they were seen during training, while non-members approach the random guessing baseline (50%), reflecting the model’s lack of prior knowledge about them.
Table 11: Membership inference evaluation with random labeling applied to both the baseline and pruned models. Forget Set is member in the baseline; pruning and RL aim to make it appear as non-member.